OpenStudy (shamallamadingdong):

The length of a rectangle is 7 1/2inches and the width is 2 3/4 inches. What is the ratio, using whole numbers, of the length to the width? 15:11 11:15 30:22 30:11

OpenStudy (shamallamadingdong):

\(\color{purple}{Pooja}\)

pooja195 (pooja195):

Write this as a improper fraction \[\huge~\rm~7\frac{1 }{ 2}\]

OpenStudy (shamallamadingdong):

\[6\frac{ 3 }{ 2 }?\]

pooja195 (pooja195):

Improper fraction...

pooja195 (pooja195):

\[\huge~\rm~Denominator \times~Unit\]

OpenStudy (shamallamadingdong):

2*6

OpenStudy (shamallamadingdong):

\[\frac{ 12 }{ 2 }\]

pooja195 (pooja195):

6?

OpenStudy (shamallamadingdong):

oh nvm

OpenStudy (shamallamadingdong):

\[\frac{ 14 }{ 2 }\]

pooja195 (pooja195):

7 x 2=14 +1=15 /2

OpenStudy (shamallamadingdong):

ok

pooja195 (pooja195):

Then we multiply the top and bottom by 2

OpenStudy (shamallamadingdong):

\[\frac{ 30 }{ 4 }\]

pooja195 (pooja195):

we get 30/4 7 1/2 = 15/2 = 30/4 2 3/4 = 11/4 length:width = (30/4):(11/4) = 30:11

OpenStudy (shamallamadingdong):

ok thx pooj

pooja195 (pooja195):

Do you understand?

OpenStudy (shamallamadingdong):

yesh

pooja195 (pooja195):

Are you sure?

OpenStudy (shamallamadingdong):

yes

pooja195 (pooja195):

okie :)